# Deciding Local Unitary Equivalence of Graph States in Quasi-Polynomial Time

### v1

@@ Abstract
We describe an algorithm with quasi-polynomial runtime $n^{\log_{2}(n)+O(1)}$ for deciding local unitary (LU) equivalence of graph states. The algorithm builds on a recent graphical characterisation of LU-equivalence via generalised local complementation. By first transforming the corresponding graphs into a standard form using usual local complementations, LU-equivalence reduces to the existence of a single generalised local complementation that maps one graph to the other. We crucially demonstrate that this reduces to solving a system of quasi-polynomially many linear equations, avoiding an exponential blow-up. As a byproduct, we generalise Bouchet’s algorithm for deciding local Clifford (LC) equivalence of graph states by allowing the addition of arbitrary linear constraints.
We also improve existing bounds on the size of graph states that are LU- but not LC-equivalent. While the smallest known examples involve 27 qubits, and it is established that no such examples exist for up to 8 qubits, we refine this bound by proving that LU- and LC-equivalence coincide for graph states involving up to 19 qubits.
@@ 2 Preliminaries
[…]
@@
where $n$ is the order of $G$ and $|G[x]|$ denotes the number of edges in the subgraph of $G$ induced by $x$ [^1] With a slight abuse of notation, $x\in\{0,1\}^{n}$ denotes the subset of vertices $\{u\in V|x_{\iota(u)}=1\}$, where $\iota:V\to[0,n-1]$ s.t. $u\prec v\Leftrightarrow\iota(u)<\iota(v)$. $\iota$ is unique as $V$ is totally ordered by $\prec$.
@@ 2.1 Generalised local complementation
We review the definition of generalised local complementation and a few of its basic properties. The reader is referred to [11] for a more detailed introduction. A generalised local complementation is a graph transformation parametrised by an independent (multi)set of vertices $S$ and a positive number $r$ called level. Like the usual local complementation, the transformation consists in toggling some of the edges of the graph depending on the number of neighbours the endpoint vertices have in common in $S$. Roughly speaking an $r$-local complementation toggles an edge if the number of common vertices in $S$ is an odd multiple of $2^{r-1}$ (an example of $2$-local complementation is given in Figure 1). To be a valid $r$-local complementation, the (multi)set $S$ on which the transformation is applied should be $r$-incident, i.e. the number of common neighbours in $S$ of any set of at most $r$ vertices should be an appropriate power of two:
@@
$u\sim_{G\star^{r}S}v\leavevmode\nobreak\ \Leftrightarrow\leavevmode\nobreak\ \left(u\sim_{G}v\leavevmode\nobreak\ \leavevmode\nobreak\ \oplus\leavevmode\nobreak\ \leavevmode\nobreak\ S\bullet\Lambda_{G}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
It is easy to see that the multiplicity in $S$ can, without loss of generality, be upperbounded by $2^{r}$.
@@
To illustrate how $r$-local complementation behaves, we consider the simple cases $r=1$ and $r=2$. First, any multiset $S$ is $1$-incident, and a 1-local complementation is nothing but a sequence of usual local complementations. 2-local complementations cannot always be decomposed into usual local complementations, it is however sufficient to consider $2$-local complementations over sets, rather than multisets (see Figure 1):
$a$$b$$c$$d$$e$$f$$a$$b$$c$$d$$e$$f$2-local complementationover $\{a,a,b,c\}$
Figure 1: Illustration on a 2-local complementation over the multiset $S=\{a,a,b,c\}$. $S$ is 2-incident: indeed $S\bullet\Lambda_{G}^{\{d,e,f\}}=2$, which is a multiple of $22^{2-1-0}=$ 2. Similarly, $S\bullet\Lambda_{G}^{\{d,e\}}=S\bullet\Lambda_{G}^{\{d,f\}}=2$ and $S\bullet\Lambda_{G}^{\{e,f\}}=4$. Edges $de$ and $df$ are toggled as $S\bullet\Lambda_{G}^{\{d,e\}}=S\bullet\Lambda_{G}^{\{d,f\}}=2\bmod 4$, but not edge $ef$ as $S\bullet\Lambda_{G}^{\{e,f\}}=0\bmod 4$. Following Proposition 2.3, the 2-local complementation over $S$ can be decomposed into a 2-local complementation over the set $\{b,c\}$ and a 1-local complementation over the set $\{a\}$.
@@
Proof 2.4.
Given a 2-local complementation over a multiset $S$, we assume without loss of generality that multiplicities are defined modulo $42^{2}=$ 4. Let $S_{1}$ be the set of vertices that have multiplicity $2$ or $3$ in $S$. Notice that $G\star^{2}S=G\star^{2}S\star^{1}S_{1}\star^{1}S_{1}$ as $S_{1}$ is an independent set, $S_{1}$ is 1-incident and generalised local complementations are self inverse. So $G\star^{2}S=G\star^{2}S\star^{2}(S_{1}\sqcup S_{1})\star^{1}S_{1}=G\star^{2}(S\sqcup S_{1}\sqcup S_{1})\star^{1}S_{1}$, where the multiplicity in the disjoint union333for any vertex $u$, $(S\sqcup S^{\prime})(u)=S(u)+S^{\prime}(u)$. $S\sqcup S_{1}\sqcup S_{1}$ is either 00 or $1$ modulo $4$. Thus the $2$-local complementation over the multiset $S$ can be decomposed into 2- and 1-local complementations over sets $\textup{{supp}}(S\sqcup S_{1}\sqcup S_{1})$ and $S_{1}$ respectively.
The 2-incidence condition can be rephrased as follows when $S$ is a set: for any subset $K$ of $V\setminus S$ of size $2$ or $3$, there is an even number of common neighbours in $S$: $|S\cap\Lambda_{G}^{K}|=0\mod 2$. In other words, the cut matrix describing the edges between $S$ and $V\setminus S$ is tri-orthogonal [3, 31, 25].
@@ 3 Algorithms for LCr- and LU-equivalences
In this section, we address the problem of deciding whether two given graphs are LU-equivalent. Additionally, we consider a variant of this problem which consists in deciding whether two graphs are LCr-equivalent for a fixed $r$. Since LU-equivalent graphs are necessarily LCr-equivalent for some $r$, the difference lies in whether the level $r$ is fixed or not.
Thanks to Proposition 2.6, if $G_{1}$ is LCr-equivalent to $G_{2}$, there exists a single $r$-local complementation, together with usual local complementations, that transforms $G_{1}$ into $G_{2}$. We introduce an algorithm that builds such a sequence of generalised local complementations, in essentially four stages:
@@
    Finally, to find such a sequence of local complementations, we apply a variant of Bouchet’s algorithm, generalised to accomodate the additional linear constraints.
Stages $(i)$, $(iii)$ and $(iv)$ can be performed in polynomial time in the order $n$ of the graphs. Stage $(ii)$ has essentially a $O(n^{r})$ time complexity, thus deciding LCr-equivalence for a fixed $r$ can be done in polynomial time. Regarding LU-equivalence, Theorem 2.5 implies $r\leq\frac{n}{2}$. We improve this upperbound and show that $r$ is at most logarithmic in $n$, leading to a quasi-polynomial time algorithm for LU-equivalence.
@@ 3.1 Bouchet algorithm, revisited
    $(A\cap D)\leavevmode\nobreak\ \Delta\leavevmode\nobreak\ (B\cap C)=V$
While the original proof involves isotropic systems [2], we provide an alternative, self-contained proof in Appendix A, that we believe to be more accessible than the original one.
@@
In this section, we consider an extension of Bouchet’s algorithm where an additional set of linear constraints on $A,B,C$ and $D$ is added as input of the problem. Such additional linear equations can reflect constraints on the applied local complementations, e.g. deciding whether two graphs are LC-equivalent under the additional constraint that all local complementations are applied on a fixed set $V_{0}$ of vertices (see Example 3.7).
While solving linear equations is easy, equation $(ii)$ is not linear, and the tour de force of Bouchet’s algorithm is to point out the fondamental properties of the solutions to both equations $(i)$ and $(ii)$ that allow to decide efficiently the LC-equivalence of graphs. In particular, Bouchet showed that a set of solutions $\mathcal{C}\subseteq\mathcal{S}$ that satisfies both $(i)$ and $(ii)$, is either small or it contains an affine subspace of $\mathcal{S}$ of small co-dimension. In the latter case, the entire set $\mathcal{S}$ is actually an affine space except for some particular cases that can be avoided by assuming that the graphs contain vertices of even degree. We extend this result as follows:
@@
Lemma 3.2 holds actually for any graph that is not in the so-called ‘Class $\alpha$’ of graphs with only odd-degree vertices together with a few additional properties444(a) any pair of non adjacent vertices should have an even number of common neighbours ; (b) for any cycle $C$, the number of triangles having a edge in $C$ is equal to the size of $C$ modulo $2$.. When the graphs are in ‘Class $\alpha$’, and in the absence of additional constraints, Bouchet proved that there is at most $2$ solutions in $\mathcal{C}$ which do not belong to the affine subspace of small codimension, and these two solutions can be easily computed (see [1], section 7). We leave as an open question the description of the set of solutions for graphs in ‘Class $\alpha$’, in particular when the two particular solutions pointed out by Bouchet do not satisfy $L$, the set of additional constraints.
From an algorithmic point of view, Lemma 3.2 leads to a straightforward generalisation of Bouchet’s algorithm to efficiently decide LC-equivalence of graphs, under a set of additional linear constraints:
@@
A Gaussian elimination can be used to compute a basis =$\mathcal{B}=\{S_{0},\ldots,S_{k}\}$ of $\mathcal{S}_{L}$, with $k<n$, in $O((n^{2}+\ell)n^{2})$ operations as there are $n^{2}$ equations in $(ii)$. If the dimension of $\mathcal{S}_{L}$ is at most $4$ (so $|\mathcal{S}_{L}|\leq 16$), we check in $O(n)$ operations, for each element of $\mathcal{S}_{L}$ whether equation $(ii)$ is satisfied. Otherwise, when $\dim(\mathcal{S}_{L})>4$, if $\mathcal{C}_{L}$ is non empty, at least one element of $\mathcal{C}_{L}$ is the sum $S_{i}+S_{j}$ of two elements of $\mathcal{B}$ (see Lemma 4.4 in [1]). For each of the $O(n^{2})$ candidates we check whether condition $(ii)$ is satisfied. If no solution is found, it implies that $\mathcal{C}_{L}=\emptyset$.
As the algorithm described in Appendix A translates a solution to (i) and (ii), into a sequence of local complementations relating two graphs, some constraints on the sequence of local complementations may be encoded as additional linear constraints. We give a fairly simple example below. A more intricate example is presented in Lemma 3.18 (in Section 3.3).
@@
An interpretation of the possible additional constraints of the extended Bouchet algorithm in terms of local Clifford operators over graph states is given in Appendix B.
@@ 3.2 Minimal local sets and standard form
We consider in this section the first stage of the LU-equivalence algorithm, which consists in putting the two input graphs into a particular shape called standard form by means of local complementations. We adapt a transformation introduced in [11], which is based on the so-called minimal local sets, and turn it into an efficient procedure.
@@
A key property of local sets is that they are invariant under LU-equivalence: Two LU-equivalent graphs $G_{1}$, $G_{2}$ have the same local sets, but not necessarily with the same generators. Moreover, the way the generators of a minimal local set differ in $G_{1}$ and $G_{2}$, provides some information on the sequence of generalised local complementations that transforms $G_{1}$ into $G_{2}$. It is thus important to cover all vertices of a graph with at least one minimal local set. Fortunately, any graph admits a minimal local set cover (MLS cover for short), and an MLS cover can be computed efficiently, within $O(n^{6.38})$ operations where $n$ is the order of the graph [10]. The information that an MLS cover provides on each vertex, is reflected by a type X, Y, Z or $\bot$, defined as follows:
@@
[…]
@@
This notion of standard form is a generalisation of the one introduced in [11], where the MLS cover considered consists of every minimal local set of the graph: $\mathcal{M}_{\text{max}}\vcentcolon=\{L\subseteq V\leavevmode\nobreak\ |\leavevmode\nobreak\ L\text{\leavevmode\nobreak\ is a minimal local set}\}$. Since there can be exponentially many minimal local sets, using $\mathcal{M}_{\text{max}}$ does not lead to an efficient procedure, for instance when computing the type of each vertex.
@@
The algorithm is fairly similar to the one presented in the proof of Proposition 24 in [11] (the main difference being that we may now add minimal local sets to the MLS cover) and can be found in Appendix C. Notice the most computationally expensive step of the algorithm is the computation of the MLS cover, hence the runtime $O(n^{6.38})$. Standard forms with respect to a common MLS cover implies some strong similarities in the structure of graphs:
@@
Lemma 3.13 was proved in [11] for the maximal MLS cover $\mathcal{M}_{\text{max}}$, but the mathematical arguments hold for any arbitrary MLS cover. A key argument is that two LU-equivalent graphs have the same vertices of type $\bot$ with respect to any arbitrary MLS cover.
After performing the algorithm described in Lemma 3.12, one can check in quadratic time [^5] In the order of the graphs, assuming the information of the types of the vertices with respect to the MLS cover is conserved. whether each vertex has the same type in $G_{1}$ and $G_{2}$, and whether every vertex of type X has the same neighbourhood in both graphs. If either condition is not met, the graphs are not LU-equivalent.
@@
If $G_{1}$ and $G_{2}$ are LCr-equivalent and are both in standard form with respect to an MLS cover $\mathcal{M}$, then $G_{1}$ and $G_{2}$ are related by a sequence of local complementations on the vertices of type $\bot$ along with a single $r$-local complementation over the vertices of type X with respect to $\mathcal{M}$.
Lemma 3.14 was proved in [11] for the maximal MLS cover $\mathcal{M}_{\text{max}}$, but the mathematical arguments hold for any arbitrary MLS cover.

### v3

@@ Abstract
We describe an algorithm with quasi-polynomial runtime $n^{\log_{2}(n)+O(1)}$ for deciding local unitary (LU) equivalence of graph states. The algorithm builds on a recent graphical characterisation of LU-equivalence via generalised local complementation. By first transforming the corresponding graphs into a standard form using usual local complementations, LU-equivalence reduces to the existence of a single generalised local complementation that maps one graph to the other. We crucially demonstrate that this reduces to solving a system of quasi-polynomially many linear equations, avoiding an exponential blow-up. As a byproduct, we generalise Bouchet’s algorithm for deciding local Clifford (LC) equivalence of graph states by allowing the addition of arbitrary linear constraints. We also improve existing bounds on the size of graph states that are LU- but not LC-equivalent. While the smallest known examples involve 27 qubits, and it is established that no such examples exist for up to 8 qubits, we refine this bound by proving that LU- and LC-equivalence coincide for graph states involving up to 19 qubits.
@@ 2 Preliminaries
Notations. Given an undirected simple graph $G=(V,E)$, we use the notation $u\sim_{G}v$ when the vertices $u$ and $v$ are connected in $G$, i.e. $(u,v)\in E$. $N_{G}(u)=\{v\in V\penalty 10000\ |\penalty 10000\ u\sim_{G}v\}$ is the neighbourhood of $u$, $Odd_{G}(D)=\{v\in V\penalty 10000\ |\penalty 10000\ |N_{G}(v)\cap D|=1\bmod 2\}$ is the odd-neighbourhood of the set $D\subseteq V$ of vertices, and $\Lambda_{G}^{D}=\{v\in V\penalty 10000\ |\penalty 10000\  \forall u\in D,u\sim_{G}v\}$ is the common neighbourhood of $D\subseteq V$. We assume $V$ totally ordered by a relation $\prec$. A local complementation with respect to a given vertex $u$ consists in complementing the subgraph induced by the neighbourhood of $u$, leading to the graph $G\star u=G\Delta K_{N_{G}(u)}$ where $\Delta$ denotes the symmetric difference on edges and $K_{A}$ is the complete graph on the vertices of $A$. With a slight abuse of notation we identify multisets of vertices with their multiplicity function $V\to\mathbb{N}$ (hence we also identify sets of vertices with their indicator functions $V\to\{0,1\}$). We consider sums of multisets: for any vertex $u$, $(S_{1}+S_{2})(u)=S_{1}(u)+S_{2}(u)$. The support $\textup{{supp}}(S)$ of a multiset $S$ of vertices denotes the set of vertices $u\in V$ such that $S(u)>0$. $S$ is said independent if no two vertices of $\textup{{supp}}(S)$ are connected. For any multiset $S$ and set $D$, $S\bullet\Lambda_{G}^{D}$ is the number of vertices of $S$, counted with their multiplicity, that are neighbours to all vertices of $D$; in other words, $S\bullet\Lambda_{G}^{D}$ is the number of common neighbours of $D$ in $S$ ($\bullet.$ is the scalar product: $A\bullet B=\sum_{u\in V}A(u).B(u)$, so $S\bullet\Lambda_{G}^{D}=\sum_{u\in\Lambda_{G}^{D}}S(u)$).
@@
where $n$ is the order of $G$ and $|G[x]|$ denotes the number of edges in the subgraph of $G$ induced by xx. [^1] With a slight abuse of notation, $x\in\{0,1\}^{n}$ denotes the subset of vertices $\{u\in V|x_{\iota(u)}=1\}$, where $\iota:V\to[0,n-1]$ s.t. $u\prec v\Leftrightarrow\iota(u)<\iota(v)$. $\iota$ is unique as $V$ is totally ordered by $\prec$.
@@ 2.1 Generalised local complementation
We review the definition of generalised local complementation and a few of its basic properties. The reader is referred to [11] for a more detailed introduction. A generalised local complementation is a graph transformation parametrised by an independent (multi)set of vertices $S$ and a positive number $r$ called level. Like the usual local complementation, the transformation consists in toggling some of the edges of the graph depending on the number of neighbours the endpoint vertices have in common in $S$. Roughly speaking an $r$-local complementation toggles an edge if the number of common vertices in $S$ is an odd multiple of $2^{r-1}$ (an example of $2$-local complementation is given in Figure˜1). To be a valid $r$-local complementation, the (multi)set $S$ on which the transformation is applied should be $r$-incident, i.e. the number of common neighbours in $S$ of any set of at most $r$ vertices should be an appropriate power of two:
@@
$u\sim_{G\star^{r}S}v\penalty 10000\ \Leftrightarrow\penalty 10000\ \left(u\sim_{G}v\penalty 10000\ \penalty 10000\ \oplus\penalty 10000\ \penalty 10000\ S\bullet\Lambda_{G}^{u,v}=2^{r-1}\bmod 2^{r}\right)$
Below we recall some basic properties of generalised local complementation.
Proposition 2.3.
[11]
  • Generalised local complementations are self inverse: if $G\star^{r}S$ is valid, then $(G\star^{r}S)\star^{r}S=G$.
  • The multiplicity in $S$ can be upperbounded by $2^{r}$: if $G\star^{r}S$ is valid, then $G\star^{r}S=G\star^{r}S^{\prime}$, where, for any vertex $u$, $S^{\prime}(u)=S(u)\bmod 2^{r}$.
  • If $G\star^{r}S_{1}$ and $G\star^{r}S_{2}$ are valid and $S_{1}+S_{2}$ is independent in $G$, then $G\star^{r}(S_{1}+S_{2})=(G\star^{r}S_{1})\star^{r}S_{2}$.
  • If $G\star^{r}S$ is valid then $G\star^{r+1}(S+S)$ is valid and induces the same transformation: $G\star^{r+1}(S+S)=G\star^{r+1}S\star^{r+1}S=G\star^{r}S$.
  • If $G\star^{r}S$ is valid then $G\star^{r-1}S$ is valid (when $r>1$) and $G\star^{r-1}S=G$.
@@
To illustrate how $r$-local complementation behaves, we consider the simple cases $r=1$ and $r=2$. First, any multiset $S$ is $1$-incident, and a 1-local complementation is nothing but a sequence of usual local complementations. 2-local complementations cannot always be decomposed into usual local complementations, it is however sufficient to consider $2$-local complementations over sets, rather than multisets (see Figure˜1):
aabbccddeeffaabbccddeeff2-local complementationover $\{a,a,b,c\}$
Figure 1: Illustration of a 2-local complementation over the multiset $\{a,a,b,c\}. SS$ is 2-incident: indeed $S\bullet\Lambda_{G}^{\{d,e,f\}}=2$, which is a multiple of $2^{2-1-0}=2$. Similarly, $S\bullet\Lambda_{G}^{\{d,e\}}=S\bullet\Lambda_{G}^{\{d,f\}}=2$ and $S\bullet\Lambda_{G}^{\{e,f\}}=4$. Edges dede and $df$ are toggled as $S\bullet\Lambda_{G}^{\{d,e\}}=S\bullet\Lambda_{G}^{\{d,f\}}=2\bmod 4$, but not edge efef as $S\bullet\Lambda_{G}^{\{e,f\}}=0\bmod 4$. Following Proposition˜2.4, the 2-local complementation over $S$ can be decomposed into a 2-local complementation over the set $\{b,c\}$ and a 1-local complementation over the set $\{a\}$.
@@
Proof 2.5.
Given a 2-local complementation over a multiset $S$, we assume without loss of generality that multiplicities are defined modulo $2^{2}=4$. Let $S_{1}$ be the set of vertices that have multiplicity $2$ or $3$ in $S$. Notice that $G\star^{2}S=G\star^{2}S\star^{1}S_{1}\star^{1}S_{1}$ as $S_{1}$ is an independent set, $S_{1}$ is 1-incident and generalised local complementations are self inverse. So $G\star^{2}S=G\star^{2}S\star^{2}(S_{1}+S_{1})\star^{1}S_{1}=G\star^{2}(S+S_{1}+S_{1})\star^{1}S_{1}$, where the multiplicity in $S+S_{1}+S_{1}$ is either 0 or $1$ modulo $4$. Thus the $2$-local complementation over the multiset $S$ can be decomposed into 2- and 1-local complementations over sets $\textup{{supp}}(S+S_{1}+S_{1})$ and $S_{1}$ respectively.
The 2-incidence condition can be rephrased as follows when $S$ is a set: for any subset $K$ of $V\setminus S$ of size 22 or $3$, there is an even number of common neighbours in $S$: $|S\cap\Lambda_{G}^{K}|=0\mod 2$. In other words, the cut matrix describing the edges between $S$ and $V\setminus S$ is tri-orthogonal [3, 31, 25].
@@ 3 Algorithms for LCr- and LU-equivalences
In this section, we address the problem of deciding whether two given graphs are LU-equivalent. Additionally, we consider a variant of this problem consisting in deciding whether two graphs are LCr-equivalent for a fixed $r$. Since LU-equivalent graphs are necessarily LCr-equivalent for some $r$, the difference lies in whether the level $r$ is fixed or not.
Thanks to Proposition˜2.7, if $G_{1}$ is LCr-equivalent to $G_{2}$, there exists a single $r$-local complementation, together with usual local complementations, that transforms $G_{1}$ into $G_{2}$. We introduce an algorithm that builds such a sequence of generalised local complementations, in essentially four stages:
@@
    Finally, to find such a sequence of local complementations, we apply a variant of Bouchet’s algorithm, generalised to accommodate the additional linear constraints.
Stages $(i)$, $(iii)$ and $(iv)$ can be performed in polynomial time in the order $n$ of the graphs. Stage $(ii)$ has essentially a $O(n^{r})$ time complexity, thus deciding LCr-equivalence for a fixed $r$ can be done in polynomial time. Regarding LU-equivalence, Theorem˜2.6 implies $r\leq\frac{n}{2}$. We improve this upperbound and show that $r$ is at most logarithmic in $n$, leading to a quasi-polynomial time algorithm for LU-equivalence.
@@ 3.1 Bouchet algorithm, revisited
    $(A\cap D)\penalty 10000\ \Delta\penalty 10000\ (B\cap C)=V$
While the original proof involves isotropic systems [2], we provide an alternative, self-contained proof in Appendix˜A, that we believe to be more accessible than the original one.
@@
In this section, we consider an extension of Bouchet’s algorithm where an additional set of linear constraints on $A,B,C$ and $D$ is added as input of the problem. Such additional linear equations can reflect constraints on the applied local complementations, e.g. deciding whether two graphs are LC-equivalent under the additional constraint that all local complementations are applied on a fixed set $V_{0}$ of vertices (see Example˜3.7).
While solving linear equations is easy, equation $(ii)$ is not linear, and the tour de force of Bouchet’s algorithm is to point out the fundamental properties of the solutions to both equations $(i)$ and $(ii)$ that allow to decide efficiently the LC-equivalence of graphs. In particular, Bouchet showed that a set of solutions $\mathcal{C}\subseteq\mathcal{S}$ that satisfies both $(i)$ and $(ii)$, is either small or it contains an affine subspace of $\mathcal{S}$ of small co-dimension. In the latter case, the entire set $\mathcal{S}$ is actually an affine space except for some particular cases that can be avoided by assuming that the graphs contain vertices of even degree. We extend this result as follows:
@@
Lemma˜3.2 holds actually for any graph that is not in the so-called ‘Class $\alpha$’ of graphs with only odd-degree vertices together with a few additional properties333(a) any pair of non adjacent vertices should have an even number of common neighbours ; (b) for any cycle $C$, the number of triangles having a edge in $C$ is equal to the size of $C$ modulo $2$.. When the graphs are in ‘Class $\alpha$’, and in the absence of additional constraints, Bouchet proved that there is at most $2$ solutions in $\mathcal{C}$ that do not belong to the affine subspace of small codimension, and these two solutions can be easily computed (see [1], section 7). We leave as an open question the description of the set of solutions for graphs in ‘Class $\alpha$’, in particular when the two particular solutions pointed out by Bouchet do not satisfy $L$, the set of additional constraints.
From an algorithmic point of view, Lemma˜3.2 leads to a straightforward generalisation of Bouchet’s algorithm to efficiently decide LC-equivalence of graphs, under a set of additional linear constraints:
@@
A Gaussian elimination can be used to compute a basis $\mathcal{B}=\{S_{0},\ldots,S_{k}\}$ of $\mathcal{S}_{L}$, with $k<n$, in $O((n^{2}+\ell)n^{2})$ operations as there are $n^{2}$ equations in $(i)$. If the dimension of $\mathcal{S}_{L}$ is at most $4$ (so $|\mathcal{S}_{L}|\leq 16$), we check in $O(n)$ operations, for each element of $\mathcal{S}_{L}$ whether equation $(ii)$ is satisfied. Otherwise, when $\dim(\mathcal{S}_{L})>4$, if $\mathcal{C}_{L}$ is non empty, at least one element of $\mathcal{C}_{L}$ is the sum $S_{i}+S_{j}$ of two elements of $\mathcal{B}$ (see Lemma 4.4 in [1]). For each of the $O(n^{2})$ candidates we check whether condition $(ii)$ is satisfied. If no solution is found, it implies that $\mathcal{C}_{L}=\emptyset$.
As the algorithm described in Appendix˜A translates a solution to (i) and (ii), into a sequence of local complementations relating two graphs, some constraints on the sequence of local complementations may be encoded as additional linear constraints. We give a fairly simple example below. A more intricate example is presented in Lemma˜3.18 (in Section˜3.3).
@@
An interpretation of the possible additional constraints of the extended Bouchet algorithm in terms of local Clifford operators over graph states is given in Appendix˜B.
@@ 3.2 Minimal local sets and standard form
We consider in this section the first stage of the LU-equivalence algorithm, consisting in putting the two input graphs into a particular shape called standard form by means of local complementations. We adapt a transformation introduced in [11], which is based on the so-called minimal local sets, and turn it into an efficient procedure.
@@
A key property of local sets is that they are invariant under LU-equivalence: Two LU-equivalent graphs $G_{1}$, $G_{2}$ have the same local sets, but not necessarily with the same generators. Moreover, the way the generators of a minimal local set differ in $G_{1}$ and $G_{2}$, provides some information on the sequence of generalised local complementations that transforms $G_{1}$ into $G_{2}$. It is thus important to cover all vertices of a graph with at least one minimal local set. Fortunately, any graph admits a minimal local set cover (MLS cover for short), and an MLS cover can be computed efficiently, within $O(n^{6.38})$ operations [^4] The algorithm presented in [10] computes a MLS cover in $O(n^{4})$ evaluations of the so-called cut-rank function, which itself can be computed in $O(n^{\omega})$ field operations where $\omega<2.38$. where $n$ is the order of the graph [10]. The information that an MLS cover provides on each vertex, is reflected by a type X, Y, Z or $\bot$, defined as follows:
@@
Given a graph $G$, a vertex $u$ is of type P $\in$ {X, Y, Z, $\bot$} with respect to a MLS cover $\mathcal{M}$, where P is
@@
This notion of standard form is a generalisation of the one introduced in [11], where the MLS cover considered consists of every minimal local set of the graph: $\mathcal{M}_{\text{max}}\vcentcolon=\{L\subseteq V\penalty 10000\ |\penalty 10000\ L\text{\penalty 10000\ is a minimal local set}\}$. Since there can be exponentially many minimal local sets, using $\mathcal{M}_{\text{max}}$ does not lead to an efficient procedure, for instance when computing the type of each vertex.
@@
The algorithm is fairly similar to the one presented in the proof of Proposition 24 in [11] (the main difference being that we may now add minimal local sets to the MLS cover) and can be found in Appendix˜C. Notice the most computationally expensive step of the algorithm is the computation of the MLS cover, hence the runtime $O(n^{6.38})$. Standard forms with respect to a common MLS cover implies some strong similarities in the structure of graphs:
@@
Lemma˜3.13 was proved in [11] for the maximal MLS cover $\mathcal{M}_{\text{max}}$, but the mathematical arguments hold for any arbitrary MLS cover. A key argument is that two LU-equivalent graphs have the same vertices of type $\bot$ with respect to any arbitrary MLS cover.
After performing the algorithm described in Lemma˜3.12, one can check in quadratic time [^5] In the order of the graphs, assuming the information of the types of the vertices with respect to the MLS cover is conserved. whether each vertex has the same type in $G_{1}$ and $G_{2}$, and whether every vertex of type X has the same neighbourhood in both graphs. If either condition is not met, the graphs are not LU-equivalent.
@@
If $G_{1}$ and $G_{2}$ are LCr-equivalent and are both in standard form with respect to an MLS cover $\mathcal{M}$, then $G_{1}$ and $G_{2}$ are related by a sequence of local complementations on the vertices of type $\bot$ along with a single $r$-local complementation over the vertices of type X.
Lemma˜3.14 was proved in [11] for the maximal MLS cover $\mathcal{M}_{\text{max}}$, but the mathematical arguments hold for any arbitrary MLS cover.